# SpatialLM: Training Large Language Models for Structured Indoor Modeling

### v1

@@ 1 Introduction
3D indoor environments are ubiquitous in our daily lives. We spend a lot of time and perform various activities in such environments every day. Therefore, a long-standing goal of artificial intelligence is to teach machines to perceive, reason about, and interact with 3D indoor scenes as humans do. In this work, we focus on the task of structured indoor modeling, which aims to extract structured indoor scene descriptions from raw sensorial inputs (i.e., RGBD scans). Specifically, the scene description includes both the architectural layout (i.e., walls, doors, and windows) and 3D object bounding boxes in the indoor environment. Such 3D structure information has been shown to benefit a wide range of real-world applications such as scene editing [40], augmented reality [4, 2], and robot navigation [17].
Compared to 3D formats such as meshes, voxels, or implicit functions, the structured scene description provides a highly compact yet flexible scene representation. In this paper, our first contribution is to regard the structured descriptions as scripts of a general-purpose language (i.e., Python) and propose to predict the language in text form, as illustrated in Figure 1. This design choice has several advantages: (i) it is human interpretable and editable; (ii) it can be easily extended to incorporate any new classes without affecting existing content in the scripts; and (iii) it allows us to leverage the strong built-in coding capability of pre-trained large language models.
@@
2 An Empirical Study on Point Cloud Feature Alignment
@@
Given the point cloud $\mathbf{P}$ (typically derived from RGBD scans), our goal is to recover the 3D layout $\mathbf{L}$, which consists of all walls, doors and windows, and a set of objects $\mathbf{O}$ in the scene. As discussed before, we represent $\mathbf{L}$ and $\mathbf{O}$ as scripts of a general-purpose language, and fine-tune LLMs to predict the language in text form. Figure 2 shows all the parameters we use for the layout and objects.
To the best of our knowledge, we are the first to employ LLMs for structured scene reconstruction from point clouds. Thus, we first propose a new dataset suitable for training the LLMs on (Section 2.1). Then, with the new dataset, we set out to investigate the following two key questions: (i) What point cloud features to use for the structured indoor modeling task (Section 2.2)? (ii) How to integrate the features into LLMs (Section 2.3)?
@@ 2.1 Dataset and Metrics
Table 1 provides a list of popular indoor RGBD scan datasets. We can make a few observations:
First, the number of scenes in real datasets is relatively small, ranging from about one hundred to a few thousands. To collect such a dataset, one needs to take RGBD scans in real scenes and manually annotate 3D information, which is a highly laborious process. Further, most real datasets provide annotations on the objects (i.e., semantic labels and 3D boxes) only. One exception is SceneCAD [1], which manually adds wall annotations to ScanNet dataset. But it only contains 1,151 rooms.
@@
ScanNet (2017) [13] real 1,513 $\bullet$
Matterport3D (2017) [7] real 90 $\bullet$
3RScan (2019) [50] real 1,482 $\bullet$
SceneCAD (2020) [1] real 1,151 $\bullet$ $\bullet$
ARKitScenes (2021) [4] real 5,048 $\bullet$
MultiScan (2022) [34] real 273 $\bullet$
ScanNet++ (2023) [59] real 1,006 $\bullet$
HM3DSem (2023) [58] real 181 $\bullet$
Structured3D (2020) [64] syn. 3,500 $\bullet$
Hypersim (2021) [45] syn. 461 $\bullet$
ProcTHOR (2022) [14] syn. 12,000 $\bullet$ $\bullet$
HSSD (2024) [14] syn. 211 $\bullet$ $\bullet$
ASE (2024) [2] syn. 100,000 $\bullet$
@@
Second, for synthetic datasets, it is much cheaper to obtain ground-truth 3D annotations. But to ensure the realism of the RGBD scans, the 3D scenes must be professionally constructed. As a result, the diversity of 3D scenes in some datasets is still limited. For example, Hypersim [45] and HSSD [24] only have 461 and 211 scenes, respectively. On the other hand, one may create a large number of scenes by employing a rule-based procedural modeling pipeline, as did in ProcTHOR [14] and ASE [2]. But the quality of the scenes is not as high.
SpatialLM dataset. In this paper, we make an effort to build a dataset suitable for training LLMs for structured indoor scene reconstruction. To this end, we take advantage of access to a large repository of interior designs from an online platform in the interior design industry. Most designs are created by professional designers and used for real-world production. We parse each 3D house into individual rooms and use a few rules to filter the rooms. This results in 12,328 distinct scenes with 54,778 rooms.
For the objects, we keep annotations on a selected set of 59 common categories (excluding wall, door, and window), and filter out small objects with side lengths all $<15\mathrm{cm}$. In the end, our dataset contains 412,932 annotated object instances of 35,426 unique CAD models.
[…]
@@
ProcTHOR [14] HSSD [24] ASE [2] SpatialLM dataset
Figure 3: Dataset visual quality comparison. The layout and object placements in ProcTHOR [14] and ASE [2] are program-generated, which exhibit noticeable differences from real-world statistics. The scenes in HSSD [24] and our dataset are fully human-authored. But HSSD only has 211 scenes.
Evaluation metrics. We evaluate our method on both layout estimation and 3D object detection. For layout estimation, let $\hat{\mathbf{L}}=\{\hat{l}_{i}\}_{i=1}^{\hat{m}}$ and $\mathbf{L}=\{l_{i}\}_{i=1}^{m}$ denote the set of predicted structural elements and ground truth, respectively. Since each element is a plane segment in 3D space, we compute the ${}_{\textrm{2D}}$) by projecting each predicted element to the plane of the ground-truth element to find the optimal assignment with Hungarian algorithm, and report the F1-scores at 0.25 and 0.5 thresholds. For 3D object detection, we match the predictions $\hat{\mathbf{O}}=\{\hat{o}_{i}\}_{i=1}^{\hat{n}}$ with the ground truth $\mathbf{O}=\{o_{i}\}_{i=1}^{n}$ via Hungarian matching. Then, we compute the ${}_{\textrm{3D}}$) between each matched pair and report the F1-scores at 0.25 and 0.5 thresholds.
@@ 2.2 Point Cloud Encoders
Extracting meaningful representations for a point cloud is challenging because of its irregular format. Learning-based approaches to process 3D point clouds can be roughly classified into three groups: Mapping-based methods are inspired by the success of 2D image backbones such as CLIP [44] and the DINO series [6, 39], and map pixel-aligned image features to 3D. For example, Lexicon3D [33] projects 3D points onto the image plane to obtain their corresponding pixel features; ConceptFusion [21] builds multimodal 3D maps via traditional SLAM and multi-view fusion approaches. Voxel-based methods first transform irregular point clouds to regular representations via 3D voxelization. Then, sparse convolution algorithms [12] can be readily used for efficient computation. Point-based methods, by contrast, process point clouds directly as sets embedded in continuous space. Recent approaches in this category [62, 54, 55, 56] have successfully explored the self-attention mechanism in Transformers and demonstrated superior performance on downstream applications.
Note that all methods discussed above produce one feature per point by default. Given that a scene-level point cloud consists of hundreds of thousands of points or more, these features are too much in terms of quantity for LLMs to consume directly. Since most modern open-source LLMs take an “Encoder-MLP-LLM” architecture for multimodal feature alignment (e.g., the Llama series [32] and Qwen series [42, 43]), we can define a point cloud encoder for LLMs as a mapping which takes a set of $N$ points as input and generates $K$ feature embeddings, where $K\ll N$:
@@
Recently, a few attempts have been made to convert point representations into feature embeddings [2, 16]. However, this step has not been carefully investigated. In this section, we conduct experiments to study (i) the effectiveness of different encoders, and (ii) the impact of $K$, i.e., the number of output features. Unless otherwise stated, Qwen2.5-0.5B [42] is employed as our base model. We use Sonata [56] as the point cloud encoder, and a two-layer MLP as the projector.
Experiment on the encoder design. We consider six possible point cloud encoders. The first group of encoders adopts a mapping-based method to fuse 2D image features to 3D points. Following Lexicon3D [33], we pool pixel-level DINOv2 features to form 3D point representations. Then, one of the two routes may be taken to reduce the points into a manageable length of tokens:
  • Voxelize ($\mathbf{P}$+DINOv2): divides the space into a fixed-resolution voxel grid and aggregates all the points in the same voxel – similar to the approach used in Scene-LLM [16];
  • Rand. sampling ($\mathbf{P}$+DINOv2): uses Farthest Point Sampling [41] to randomly sample a fixed number of points with their features.
From Table 2 we can see, these two options result in extremely low F1-scores. We hypothesize that this is because too much spatial information is lost during the naive down-sampling step. While this may be acceptable for certain applications such as visual Q&A, it is problematic for our task which relies heavily on the spatial information to reconstruct the geometric coordinates of objects.
@@
Encoder F1 - layout F1 - object
@@
Voxelize ($\mathbf{P}$+DINOv2) 10.7 5.6 0.1 0.0
Rand. sampling ($\mathbf{P}$+DINOv2) 9.7 5.5 0.2 0.0
3DCNN enc. ($\mathbf{P}$) 75.2 71.6 59.8 46.1
3DCNN enc. ($\mathbf{P}$+DINOv2) 77.4 75.0 57.1 45.0
3DCNN enc. ($\mathbf{P}$) + Voxelize (DINOv2) 79.4 77.2 62.9 46.7
Sonata/PTv3 enc. ($\mathbf{P}$) 80.5 77.6 65.1 49.4
The second group of encoders adopts a standard 3DCNN encoder to generate pooled features. We follow SceneScript [2] to build our encoder with a sparse 3D convolution library [48]. Here we consider three variants:
@@
As shown in Table 2, a learned encoder is effective in keeping the necessary semantic and geometric information for reconstruction. We also find that incorporating DINOv2 features yields slightly better results, possibly due to the enhanced contextual information.
The last encoder we consider is Sonata [56], a variant of Point Transformer V3 (PTv3) [55] which removes the decoder and focuses on self-supervised learning on the encoder. Compared to other backbones, such an “encoder-only” design makes it quite convenient for adapting to LLMs. Indeed, as shown in Table 2, with raw points as input, it already achieves the best results among all encoders.
@@
1$\times$ 123 80.5 77.6 65.1 49.4
1.25$\times$ 197 80.7 78.4 65.3 52.2
1.5$\times$ 286 82.9 80.7 68.7 58.6
2$\times$ 510 83.4 81.4 71.6 61.1
2.5$\times$ 788 82.4 79.4 72.2 61.7
4$\times$ 2,000 81.4 79.3 69.9 58.9
Experiment on the spatial resolution. Recall in Eq. 1 that the parameter $K$ denotes the number of feature embeddings or visual tokens, which is critical to the performance of LLMs. In this experiment, we study the effect of $K$.
However, fixing $K$ is difficult because different point clouds generally produce varying $K$ values. Instead, since the encoders employ a coarse-to-fine structure for point cloud processing, we set the resolution at the finest level at $5\mathrm{cm}$ and use a 5-level hierarchical structure by default. Then, we gradually increase the resolution up to 4$\times$, and report the results in Table 3. As can be seen, model performance steadily improves from 1$\times$ to 2$\times$ resolutions, but further increasing the resolution hurts the performance. This is likely due to the excessively long token sequences.
2.3 Training Schedules
Training stages F1 - layout F1 - object

### v2

@@ 1 Introduction
3D indoor environments are ubiquitous in our daily lives. We spend a lot of time and perform various activities in such environments every day. Therefore, a long-standing goal of artificial intelligence is to teach machines to perceive, reason about, and interact with 3D indoor scenes as humans do. In this work, we focus on the task of structured indoor modeling, which aims to extract structured indoor scene descriptions from raw sensorial inputs (i.e., RGBD scans). Specifically, the scene description includes both the architectural layout (i.e., walls, doors, and windows) and 3D object bounding boxes in the indoor environment. Such 3D structure information has been shown to benefit a wide range of real-world applications such as scene editing [PatilPLFSZ24], augmented reality [ARKitScenes, SceneScript], and robot navigation [GuSWTW22].
Compared to 3D formats such as meshes, voxels, or implicit functions, the structured scene description provides a highly compact yet flexible scene representation. In this paper, our first contribution is to regard the structured descriptions as scripts of a general-purpose language (i.e., Python) and propose to predict the language in text form, as illustrated in Figure˜1. This design choice has several advantages: (i) it is human interpretable and editable; (ii) it can be easily extended to incorporate any new classes without affecting existing content in the scripts; and (iii) it allows us to leverage the strong built-in coding capability of pre-trained large language models.
@@
2 Related Work
2.1 Structured Indoor Modeling
Structured indoor modeling aims to recover structural elements (i.e., walls, doors, and windows) as well as 3D objects from unstructured point clouds. In the literature, some studies focus on room layout estimation from point clouds. Early methods [MuraliSOP17, abs-1907-00631] rely on traditional geometric analysis algorithms (i.e., RANSAC) to detect wall planes. SceneCAD [AvetisyanKCDDN20] jointly optimizes the layout estimation and CAD model alignment of the scanned scene. RoomFormer [YueKSE23] proposes to train a Transformer-based network to predict corners and rooms. Meanwhile, a number of recent methods [MisraGJ21, RukhovichV022, WangDDSLLL022, 0001GYLLW00G24, KolodiazhnyiVSR25] have advanced the state-of-the-art on 3D object detection from point clouds.
The closest work to ours is SceneScript [SceneScript], which also casts structured indoor reconstruction as a sequence modeling problem. It demonstrates great flexibility as the language can be tailored to different scenarios without changing the method. But unlike our method, SceneScript uses domain-specific tokens for language commands, thus requiring a specialized decoder. Instead, SpatialLM strictly adheres to the “Encoder-MLP-LLM” architecture and is directly fine-tuned based on modern foundation models.
2.2 Large Language Models for 3D Scene Understanding
Recently, there has been a growing interest in empowering LLMs to understand and reason about 3D spaces, as defined in public benchmarks such as ScanRefer [ChenCN20] and ScanQA [AzumaMKK22]. One line of work [Chat-3D, HuangYMLLW0ZJ024, Chat-Scene, Uni3D-LLM, SceneVerse] first uses 3D detectors to locate objects in the point cloud, then extracts features of individual objects using a pre-trained object-level encoder. In this paper, we study how such 3D detectors can be learned with modern LLMs.
Other methods directly extract scene-level features from the point cloud and align the features with LLMs [HongZCZDCG23, ZhuWZCL24, ChenCZLY0Z0024, Scene-LLM]. 3D-LLM [HongZCZDCG23] uses pre-trained image encoder to obtain 2D features from multi-view images and back-projects onto the point cloud [JatavallabhulaK23]. It then leverages a Q-former [BLIP2] to map them to a fixed number of tokens as input to the LLM. LL3DA [ChenCZLY0Z0024] adopts a 3D visual encoder pre-trained on ScanNet detection as the scene encoder. ScanReason [ZhuWZCL24] further employs a 3D point cloud encoder to obtain fine-grained geometric information for accurate object grounding. Scene-LLM [Scene-LLM] tokenizes the 3D point cloud features by simply dividing the space into a fixed-resolution voxel grid and aggregating features in the same voxel.
Besides point cloud-based representations, recent studies have explored other ways to integrate 3D information in LLMs. SpatialVLM [0003XKISGX24] primarily focuses on synthesizing large-scale 3D VQA data to enhance the spatial reasoning capabilities of the VLM. LLaVA-3D [LLaVA-3D] injects 3D position embeddings into 2D CLIP features to establish a unified architecture for 2D image understanding and 3D scene understanding. SpatialRGPT [ChengYFGYK0L24] extracts a 3D scene graph from a single image to enhance grounded spatial reasoning in LLMs. Video-3D LLM [Video-3D-LLM] learns representations from video frames and 3D coordinates to leverage the strong priors of pre-trained 2D Video LLMs. LSceneLLM [LSceneLLM] uses an adaptive self-attention module to capture fine-grained details in local areas. Our work differs from these studies in that we focus on analyzing the structural information of the scene.
3 SpatialLM
@@
Given the point cloud $\mathbf{P}$ (typically derived from RGBD scans), our goal is to recover the 3D layout $\mathbf{L}$, which consists of all walls, doors and windows, and a set of objects $\mathbf{O}$ in the scene. As discussed before, we represent $\mathbf{L}$ and $\mathbf{O}$ as scripts of a general-purpose language, and fine-tune LLMs to predict the language in text form. Figure˜2 shows all the parameters we use for the layout and objects.
To the best of our knowledge, we are the first to employ LLMs for structured scene reconstruction from point clouds. Thus, we first propose a new dataset suitable for training the LLMs on (Section˜3.1). Then, with the new dataset, we set out to investigate the following two key questions: (i) What point cloud features to use for the structured indoor modeling task (Section˜3.2)? (ii) How to integrate the features into LLMs (Section˜3.3)?
@@ 3.1 Dataset and Metrics
ScanNet (2017) [ScanNet] real 1,513 $\bullet$
Matterport3D (2017) [Matterport3D] real 90 $\bullet$
3RScan (2019) [WaldANTN19] real 1,482 $\bullet$
SceneCAD (2020) [AvetisyanKCDDN20] real 1,151 $\bullet$ $\bullet$
ARKitScenes (2021) [ARKitScenes] real 5,047 $\bullet$
MultiScan (2022) [MaoZJCS22] real 273 $\bullet$
ScanNet++ (2023) [ScanNetpp] real 1,006 $\bullet$
HM3DSem (2023) [YadavRRGTGMCBSC23] real 181 $\bullet$
Structured3D (2020) [Structured3D] syn. 3,500 $\bullet$
Hypersim (2021) [Hypersim] syn. 461 $\bullet$
ProcTHOR (2022) [ProcTHOR] syn. 12,000 $\bullet$ $\bullet$
HSSD (2024) [ProcTHOR] syn. 211 $\bullet$ $\bullet$
ASE (2024) [SceneScript] syn. 100,000 $\bullet$
@@
Table˜1 provides a list of popular indoor RGBD scan datasets. We can make a few observations:
First, the number of scenes in real datasets is relatively small, ranging from about one hundred to a few thousands. To collect such a dataset, one needs to take RGBD scans in real scenes and manually annotate 3D information, which is a highly laborious process. Further, most real datasets provide annotations on the objects (i.e., semantic labels and 3D boxes) only. One exception is SceneCAD [AvetisyanKCDDN20], which manually adds wall annotations to ScanNet dataset. But it only contains 1,151 rooms.
Second, for synthetic datasets, it is much cheaper to obtain ground-truth 3D annotations. But to ensure the realism of the RGBD scans, the 3D scenes must be professionally constructed. As a result, the diversity of 3D scenes in some datasets is still limited. For example, Hypersim [Hypersim] and HSSD [HSSD] only have 461 and 211 scenes, respectively. On the other hand, one may create a large number of scenes by employing a rule-based procedural modeling pipeline, as did in ProcTHOR [ProcTHOR] and ASE [SceneScript]. But the quality of the scenes is not as high.
SpatialLM dataset. In this paper, we make an effort to build a dataset suitable for training LLMs for structured indoor scene reconstruction. To this end, we take advantage of access to a large repository of interior designs from an online platform [^1] https://www.kujiale.com/ in the interior design industry. Most designs are created by professional designers and used for real-world production. We parse each 3D house into individual rooms and use a few rules to filter the rooms. This results in 12,328 distinct scenes with 54,778 rooms.
For the objects, we keep annotations on a selected set of 59 common categories (excluding wall, door, and window), and filter out small objects with side lengths all $15\mathrm{cm}$. In the end, our dataset contains 412,932 annotated object instances of 35,426 unique CAD models.
For each scene, we generate photo-realistic RGBD images utilizing an industry-leading rendering engine. At the time of rendering, a camera trajectory traversing each room is simulated, based on which the images are taken at an interval of $0.5\,\textrm{m}$. Figure˜3 compares the visual quality of our dataset to several recent synthetic datasets. Finally, the dataset is divided into 11,328/500/500 scenes for training/validation/testing.
@@
ProcTHOR [ProcTHOR] HSSD [HSSD] ASE [SceneScript] SpatialLM dataset
Figure 3: Dataset visual quality comparison. The layout and object placements in ProcTHOR [ProcTHOR] and ASE [SceneScript] are program-generated, which exhibit noticeable differences from real-world statistics. The scenes in HSSD [HSSD] and our dataset are fully human-authored. But HSSD only has 211 scenes.
[…]
@@ 3.2 Point Cloud Encoders
Extracting meaningful representations for a point cloud is challenging because of its irregular format. Learning-based approaches to process 3D point clouds can be roughly classified into three groups: Mapping-based methods are inspired by the success of 2D image backbones such as CLIP [CLIP] and the DINO series [DINO, DINOv2], and map pixel-aligned image features to 3D. For example, Lexicon3D [Lexicon3D] projects 3D points onto the image plane to obtain their corresponding pixel features; ConceptFusion [JatavallabhulaK23] builds multimodal 3D maps via traditional SLAM and multi-view fusion approaches. Voxel-based methods first transform irregular point clouds to regular representations via 3D voxelization. Then, sparse convolution algorithms [ChoyGS19] can be readily used for efficient computation. Point-based methods, by contrast, process point clouds directly as sets embedded in continuous space. Recent approaches in this category [ZhaoJJTK21, 0002LJLZ22, PTv3, Sonata] have successfully explored the self-attention mechanism in Transformers and demonstrated superior performance on downstream applications.
Note that all methods discussed above produce one feature per point by default. Given that a scene-level point cloud consists of hundreds of thousands of points or more, these features are too much in terms of quantity for LLMs to consume directly. Since most modern open-source LLMs take an “Encoder-MLP-LLM” architecture for multimodal feature alignment (e.g., the Llama series [llama3] and Qwen series [qwen2.5, qwen2.5-vl]), we can define a point cloud encoder for LLMs as a mapping which takes a set of $N$ points as input and generates $K$ feature embeddings, where $K\ll N$:
@@
Recently, a few attempts have been made to convert point representations into feature embeddings [SceneScript, Scene-LLM]. However, this step has not been carefully investigated. In this section, we conduct experiments to study (i) the effectiveness of different encoders, and (ii) the impact of $K$, i.e., the number of output features. Unless otherwise stated, Qwen2.5-0.5B [qwen2.5] is employed as our base model. We use Sonata [Sonata] as the point cloud encoder, and a two-layer MLP as the projector.
Experiment on the encoder design. We consider six possible point cloud encoders. The first group of encoders adopts a mapping-based method to fuse 2D image features to 3D points. Following Lexicon3D [Lexicon3D], we pool pixel-level DINOv2 features to form 3D point representations. Then, one of the two routes may be taken to reduce the points into a manageable length of tokens:
  • Voxelize ($\mathbf{P}$+DINOv2): divides the space into a fixed-resolution voxel grid and aggregates all the points in the same voxel – similar to the approach used in Scene-LLM [Scene-LLM];
  • Rand. sampling ($\mathbf{P}$+DINOv2): uses Farthest Point Sampling [QiYSG17] to randomly sample a fixed number of points with their features.
From Table˜2 we can see, these two options result in extremely low F1-scores. We hypothesize that this is because too much spatial information is lost during the naive down-sampling step. While this may be acceptable for certain applications, such as visual Q&A, it is problematic for our task, which relies heavily on the spatial information to reconstruct the geometric coordinates of objects.
@@
Encoder Params F1 - layout F1 - object
@@
Voxelize ($\mathbf{P}$+DINOv2) 0.0 10.8 5.5 0.1 0.0
Rand. sampling ($\mathbf{P}$+DINOv2) 0.0 10.6 4.9 0.2 0.0
3DCNN enc. ($\mathbf{P}$) 10.5 79.4 76.3 59.8 46.1
3DCNN enc. ($\mathbf{P}$+DINOv2) 11.6 81.5 79.4 57.1 45.0
3DCNN enc. ($\mathbf{P}$) + Voxelize (DINOv2) 11.9 83.8 81.6 62.9 46.7
Sonata/PTv3 enc. ($\mathbf{P}$) 108.8 84.6 82.6 65.1 49.4
The second group of encoders adopts a standard 3DCNN encoder to generate pooled features. We follow SceneScript [SceneScript] to build our encoder with a sparse 3D convolution library [TangLL0022]. Here we consider three variants:
@@
As shown in Table˜2, a learned encoder is effective in keeping the necessary semantic and geometric information for reconstruction. We also find that incorporating DINOv2 features yields slightly better results, possibly due to the enhanced contextual information.
The last encoder we consider is Sonata [Sonata], a variant of Point Transformer V3 (PTv3) [PTv3] which removes the decoder and focuses on self-supervised learning on the encoder. Compared to other backbones, such an “encoder-only” design makes it quite convenient for adapting to LLMs. Indeed, as shown in Table˜2, with raw points as input, it already achieves the best results among all encoders.
@@
$\times 1$23 84.6 82.6 65.1 49.4
@@
Experiment on the spatial resolution. Recall in Eq.˜1 that the parameter $K$ denotes the number of feature embeddings or visual tokens, which is critical to the performance of LLMs. In this experiment, we study the effect of $K$.
@@
3.3 Training Schedules
Training stages F1 - layout F1 - object